An ion source produces a beam of ions, each with charge +8 e and mass 2.505e-26 kg. The ions are accelerated by a potential difference of 33 kV and pass horizontally into a region where there is a uniform magnetic field of magnitude BB = 1.8 tesla. Calculate the strength of the smallest electric field, to be set up in the same region, that will allow the ions to pass through undeflected.
Answer #1

using conservation of energy

qV = 0.5 m v^2

8* 1.6* 10^-19* 33000 = 0.5* 2.505* 10^-26* v^2

v = 1.836* 10^6 m/s

now as we also know

E = B v

E = 1.8*1.836* 10^6

E = 3.306 * 10^6 N/C

=======

Comment before rate in case any doubt, will reply for sure.. goodluck